# Unified Triplet-Level Hallucination Evaluation for Large Vision-Language Models

### v2

@@ E.1 Prompt for triplets extraction with GPT-4
E.2 Prompt for question generation with GPT-4V
@@
E.3 Prompt for GPT-4 Judge
The prompt for our proposed GPT-4 judge method is illustrated in Figure 8.
Figure 8: Prompt for the GPT-4 Judge method.
E.4 Prompt for LVLMs in Evaluation
The prompt for generating responses from LVLMs for evaluation is the combination of a question and the corresponding image.
Appendix F Human Annotation Guideline
@@
Appendix G More Discussion
@@
Currently, the proposed triplet-level evaluation is primarily deployed on LVLMs, whose extension to diffusion models [43, 44, 45, 46, 47, 48] is feasible, while for the hallucination mitigation proposed in §5.4 can be further enhanced by utilizing stronger vision encoder [49, 50, 51, 52] and visual tools (e.g., object detectors [53, 54]) to better extract visual information for LLM reasoning.

### v3

@@ A.1 Prompt for triplets extraction with GPT-4
A.2 Prompt for LLM Judge
The prompt for our proposed LLM judge method is illustrated in Figure 6.
Figure 6: Prompt for the LLM Judge method.
A.3 Prompt for question generation with GPT-4V
@@
A.4 Prompts for Evaluating LVLMs
When evaluating LVLMs on Tri-HE, the prompt we use is the question itself. Questions are fed into LVLMs along with the corresponding images.
Appendix B Configurations for LVLM Evaluation
For LVLM evaluations, we directly use the default configuration settings provided in their publicly available code repositories. For instance, the configurations utilized for evaluating LLaVA models are accessible at https://github.com/haotian-liu/LLaVA.
Appendix C Human Annotation Guideline
@@
Appendix D More Discussion
@@
Currently, the proposed triplet-level evaluation is primarily deployed on LVLMs, whose extension to diffusion models (Chen et al., 2023c; Gao et al., 2023; 2024a; 2024b; Li et al., 2023d; Liu et al., 2023c; Wang et al., 2024) is feasible, while for the hallucination mitigation proposed in §5.4 can be further enhanced by utilizing stronger vision encoder (Chen et al., 2021; 2023a; Liu et al., 2022; Zhili et al., 2023) and visual tools (e.g., object detectors (Han et al., 2021; Li et al., 2022)) to better extract visual information for LLM reasoning.